OpenStudy (lissiewalkin):

I am asked to find the mole fraction of SCN- The information given is that a test tube contains 4ml of 0.0020M Fe(NO3)3 + and 6ml of 0.0020M KSCN. How would I calculate this? Please help!

OpenStudy (abb0t):

First, you need to find the moles of each, which is why they gave you volume. Remember that M = \(\sf \frac{mol}{L}\) Uuse that to cancel out the units for liter Also, do not forget to change mL \(\rightarrow\) L units.

OpenStudy (lissiewalkin):

I followed what you said about finding the moles. I converted 4ml to liters and multiplied that by 0.0020mol/L. And I got 0.000008mol Fe(NO3)3. Is that correct?

OpenStudy (abb0t):

Yes, now to find the mole fraction of each one. It should be equal to ONE. Meaning \(\sf \frac{moles~of~A}{total~moles~of~A+B}\) DO THAT FOR EACH

OpenStudy (abb0t):

So \(\sf \frac{0.000008~mol}{0.000008~mol~A + n~mol~of~B}\)

OpenStudy (lissiewalkin):

Ohh its starting to all make sense now. I got (0.000012mol KSCN/ 0.000008mol + 0.000012mol) = 0.6 Would that be the mole fraction? Because I'm asked to find the mole fraction of SCN- but the calculations have KSCN-. Do I just ignore the K?

OpenStudy (joannablackwelder):

I think you can for this problem. Because there is 1 mol SCN- for each mol of KSCN.

OpenStudy (lissiewalkin):

Oh ok that makes sense! Thanks @JoannaBlackwelder and @abb0t! I really appreciate it :]